# On the solution of the Collatz problem

### v26

@@ 3.2. Properties of Collatz’s Sequences
and by the formula $03A_{1,k-1}-2A_{1,k}=$ 0 there are
$A_{1,k}=3A_{1,k-1}/2=2^{\eta-k+1+\sigma(s)}3^{k-2}\texttt{ for }k=3,4,\ldots,\eta_{0}.$
So we obtain Formula 3.6.
For $\xi=0,1,2,\ldots,s-2$, if $k=\eta_{\xi}+1$ then by the formula $03A_{1,k-1}-2^{1+r_{\xi+1}}A_{1,k}=$ 0, there are
$A_{1,k}=3A_{k-1,1}/2^{1+r_{\xi+1}}=2^{\eta-k+1+\sigma(\xi+2,s)}3^{k-2},$
else by the formula $03A_{1,k-1}-2A_{1,k}=$ 0 there are
$A_{1,k}=3A_{1,k-1}/2=2^{\eta-k+1+\sigma(\xi+2,s)}3^{k-2}\texttt{ for }k=\eta_{\xi}+2,\eta_{\xi}+3,\ldots,\eta_{\xi+1}.$
So we obtain Formula 3.7.
Thus by formulas 3.6-3.7 the first collum of the matrix $\mathbf{A}^{*}$ is determined.
@@
and by the formula $03A_{1,k}-2^{1+r_{s}}A_{\eta,k}=$ 0 there are
$A_{\eta,k}=3A_{1,k}/2^{1+r_{s}}=2^{\eta-k+\sigma(s-1)}3^{k-1}\texttt{ for }k=2,3,\ldots,\eta_{0}.$
So we obtain Formula 3.8.
For $\xi=0,1,2,\ldots,s-2$, if $k=\eta_{\xi}+1$ then by the formula $03A_{1,k}-2^{1+r_{s}}A_{\eta,k}=$ 0 there are
$A_{\eta,k}=3A_{1,k}/2^{1+r_{s}}=2^{\eta-k+\sigma(\xi+2,s-1)}3^{k-1},$
else by the formula $03A_{1,k}-2^{1+r_{s}}A_{\eta,k}=$ 0, there are
$A_{\eta,k}=3A_{1,k}/2^{1+r_{s}}=2^{\eta-k+\sigma(\xi+2,s-1)}3^{k-1}\texttt{ for }k=\eta_{\xi}+2,\eta_{\xi}+3,\ldots,\eta_{\xi+1}.$
So we obtain Formula 3.9.
Thus by formulas 3.8-3.9 the last collum of the matrix $\mathbf{A}^{*}$ is determined.
Second, we determine the first row of the matrix $\mathbf{A}^{*}$. By the formula
$3A_{k,1}-2A_{k-1,1}=0\texttt{ for }k=2,3,\ldots,\eta_{0},$
we obtain Formula 3.10.
For $\xi=0,1,2,\ldots,s-2$, if $k=\eta_{\xi}+1$ then by the formula $03A_{k,1}-2^{1+r_{\xi+1}}A_{k-1,1}=$ 0 there are
$A_{k,1}=2^{1+r_{\xi+1}}A_{k-1,1}/3=2^{k-1+\sigma(\xi+1)}3^{\eta-k},$
else by the formula $03A_{k,1}-2A_{k-1,1}=$ 0, if $\xi<s-2$ then there are
$A_{k,1}=2A_{k-1,1}/3=2^{k-1+\sigma(\xi+1)}3^{\eta-k}\texttt{ for }k=\eta_{\xi}+2,\eta_{\xi}+3,\ldots,\eta_{\xi+1}.$
So we obtain Formula 3.11.
Thus by formulas 3.10-3.11 the first row of the matrix $\mathbf{A}^{*}$ is determined.
Third, we determine the $\kappa^{th}$ row of the matrix $\mathbf{A}^{*}$ where $\kappa=3,4,\ldots,\eta$ and positive integers $A_{k,\kappa}$ for $k=2,3,\ldots,\kappa-1$. By the formula $03A_{k,\kappa}-2A_{k-1,\kappa}=$ 0 we obtain Formula 3.12.
If $\kappa-1>\eta_{0}$ then let $\eta_{\tau-1}<\kappa-1\leq\eta_{\tau}$. Then for $\xi=0,1,2,\ldots,\tau-1$, if $k=\eta_{\xi}+1$ then by the formula $03A_{k,\kappa}-2^{1+r_{\xi+1}}A_{k-1,\kappa}=$ 0 there are
$A_{k,\kappa}=2^{1+r_{\xi+1}}A_{k-1,\kappa}/3=2^{k-1+\sigma(\xi+1)}3^{\eta-k},$
else by the formula $03A_{k,\kappa}-2A_{k-1,\kappa}=$ 0 there are
$A_{k,\kappa}=2A_{k-1,\kappa}/3=2^{k-1+\sigma(\xi+1)}3^{\eta-k}$
for $k=\eta_{\xi}+2,\eta_{\xi}+3,\ldots,\min(\eta_{\xi+1},\kappa-1)$. So we obtain Formula 3.13.
Thus by formulas 3.12-3.13 integers $A_{k,\kappa}$ are determined for $\kappa=3,4,\ldots,\eta$ and $k=2,3,\ldots,\kappa-1$.
@@
Finally, we determine the $\kappa^{th}$ row of the matrix $\mathbf{A}^{*}$ where $\kappa=2,3,\ldots,\eta-1$ and positive integers $A_{\kappa+k,\kappa}$ for $k=1,2,\ldots,\eta-\kappa$. By the formula
$03A_{\kappa+k,\kappa}-2A_{\kappa+k-1,\kappa}=$ 0
we obtain Formula 3.14.
If $\eta-\kappa>\eta_{0}$ then let $\eta_{\tau-1}<\eta-\kappa\leq\eta_{\tau}$. Then for $\xi=0,1,2,\ldots,\tau-1$, if $k=\eta_{\xi}+1$ then by the formula $03A_{\kappa+k,\kappa}-2^{1+r_{\xi+1}}A_{\kappa+k-1,\kappa}=$ 0 there are
$A_{\kappa+k,\kappa}=2^{1+r_{\xi+1}}A_{\kappa+k-1,\kappa}/3=2^{k+\sigma(\xi+1)}3^{\eta-1-k},$
else by the formula $03A_{\kappa+k,\kappa}-2A_{\kappa+k-1,\kappa}=$ 0 there are
$A_{\kappa+k,\kappa}=2A_{\kappa+k-1,\kappa}/3=2^{k+\sigma(\xi+1)}3^{\eta-1-k}$
for $k=\eta_{\xi}+2,\eta_{\xi}+3,\ldots,\min(\eta_{\xi+1},\eta_{\xi}+\eta-\kappa)$. So we obtain Formula 3.15.
Thus by formulas 3.14-3.15 integers $A_{\kappa+k,\kappa}$ are determined for $\kappa=2,3,\ldots,\eta-1$ and $k=1,2,\ldots,\eta-\kappa$.
Hence the matrix $\mathbf{A}^{*}$ is determined as Formula 3.5.

### v32

@@ 3.2. Properties of Collatz’s Sequences
and by the formula $3A_{1,\kappa-1}-2A_{1,\kappa}=0$ for $\kappa=3,4,\ldots,\eta_{1}$ there are
$A_{1,\kappa}=3A_{1,\kappa-1}/2=2^{\eta-\kappa+1+\sigma(s)}3^{\kappa-2}\texttt{ for }\kappa=3,4,\ldots,\eta_{1}.$
So we obtain
(3.8) $A_{1,\kappa}=2^{\eta-\kappa+1+\sigma(s)}3^{\kappa-2}\texttt{ for }\kappa=2,3,\ldots,\eta_{1}.$
For $\xi=1,2,\ldots,s-1$ and $\kappa=\eta_{\xi}+1,\eta_{\xi}+2,\ldots,\eta_{\xi+1}$, if $\kappa=\eta_{\xi}+1$ then by the formula $3A_{1,\kappa-1}-2^{1+r_{\xi}}A_{1,\kappa}=0$, there are
$A_{1,\kappa}=3A_{1,\kappa-1}/2^{1+r_{\xi}}=2^{\eta-\kappa+1+\sigma(\xi+1,s)}3^{\kappa-2},$
else by the formula $3A_{1,\kappa-1}-2A_{1,\kappa}=0$ for $\kappa=\eta_{\xi}+2,\eta_{\xi}+3,\ldots,\eta_{\xi+1}$ there are
$A_{1,\kappa}=3A_{1,\kappa-1}/2=2^{\eta-\kappa+1+\sigma(\xi+1,s)}3^{\kappa-2}\texttt{ for }\kappa=\eta_{\xi}+2,\eta_{\xi}+3,\ldots,\eta_{\xi+1}.$
So for $\xi=1,2,\ldots,s-1$ we obtain
(3.9) $A_{1,\kappa}=2^{\eta-\kappa+1+\sigma(\xi+1,s)}3^{\kappa-2}\texttt{ for }\kappa=\eta_{\xi}+1,\eta_{\xi}+2,\ldots,\eta_{\xi+1}.$
Thus by formulas 3.8-3.9 the first collum of the matrix $\mathbf{A}^{*}$ is determined.
@@
and by the formula $3A_{1,\kappa}-2^{1+r_{s}}A_{\eta,\kappa}=0$ for $\kappa=2,3,\ldots,\eta_{1}$ there are
$A_{\eta,\kappa}=3A_{1,\kappa}/2^{1+r_{s}}=2^{\eta-\kappa+\sigma(s-1)}3^{\kappa-1}\texttt{ for }\kappa=2,3,\ldots,\eta_{1}.$
So we obtain
(3.10) $A_{\eta,\kappa}=2^{\eta-\kappa+\sigma(s-1)}3^{\kappa-1}\texttt{ for }\kappa=1,2,\ldots,\eta_{1}.$
For $\xi=1,2,\ldots,s-1$, by the formula $3A_{1,\kappa}-2^{1+r_{s}}A_{\eta,\kappa}=0$ there are
$A_{\eta,\kappa}=3A_{1,\kappa}/2^{1+r_{s}}=2^{\eta-\kappa+\sigma(\xi+1,s-1)}3^{\kappa-1}$
for $\kappa=\eta_{\xi}+1,\eta_{\xi}+2,\ldots,\min(\eta_{\xi+1},\eta-1)$. So for $\xi=1,2,\ldots,s-1$ we obtain
(3.11) $A_{\eta,\kappa}=2^{\eta-\kappa+\sigma(\xi+1,s-1)}3^{\kappa-1}\texttt{ for }\kappa=\eta_{\xi}+1,\eta_{\xi}+2,\ldots,\min(\eta_{\xi+1},\eta-1).$
Thus by formulas 3.10-3.11 the last collum of the matrix $\mathbf{A}^{*}$ is determined.
@@
Second, we determine the $\kappa^{th}$ row of the matrix $\mathbf{A}^{*}$ where $\kappa=1,2,\ldots,\eta-1$ and positive integers $A_{k,\kappa}$ for $k=\kappa+1,\kappa+2,\ldots,\eta$.
For $\xi=0,1,2,\ldots,s-1$ and $\kappa=\eta_{\xi}+1,\eta_{\xi}+2,\ldots,\min(\eta_{\xi+1},\eta-1)$, by the formula $3A_{k,\kappa}-2A_{k-1,\kappa}=0$ for $k=\kappa+1,\kappa+2,\ldots,\eta_{\xi+1}$ there are
$A_{k,\kappa}=2A_{k-1,\kappa}/3=2^{k-\kappa}3^{\eta-k+\kappa-1}\texttt{ for }k=\kappa+1,\kappa+2,\ldots,\eta_{\xi+1},$
so that we obtain
(3.12) $A_{k,\kappa}=2^{k-\kappa}3^{\eta-k+\kappa-1}\texttt{ for }k=\kappa+1,\kappa+2,\ldots,\eta_{\xi+1},$
and for $\zeta=\xi+1,\xi+2,\ldots,s-1$ and $k=\eta_{\zeta}+1,\eta_{\zeta}+2,\ldots,\eta_{\zeta+1}$, if $k=\eta_{\zeta}+1$ then by the formula $3A_{k,\kappa}-2^{1+r_{\zeta+1}}A_{k-1,\kappa}=0$ there are
$A_{k,\kappa}=2^{1+r_{\zeta+1}}A_{k-1,\kappa}/3=2^{k-\kappa+\sigma(\xi+1,\zeta)}3^{\eta-k+\kappa-1},$
else by the formula $3A_{k,\kappa}-2A_{k-1,\kappa}=0$ there are
$A_{k,\kappa}=2A_{k-1,\kappa}/3=2^{k-\kappa+\sigma(\xi+1,\zeta)}3^{\eta-k+\kappa-1}\texttt{ for }k=\eta_{\zeta}+2,\eta_{\zeta}+3,\ldots,\eta_{\zeta+1},$
so that for $\zeta=\xi+1,\xi+2,\ldots,s-1$ we obtain
(3.13) $A_{k,\kappa}=2^{k-\kappa+\sigma(\xi+1,\zeta)}3^{\eta-k+\kappa-1}\texttt{ for }k=\eta_{\zeta}+1,\eta_{\zeta}+2,\ldots,\eta_{\zeta+1}.$
Thus by formulas 3.12-3.13 integers $A_{k,\kappa}$ are determined for $\kappa=1,2,\ldots,\eta-1$ and $k=\kappa+1,\kappa+2,\ldots,\eta$.
Finally, we determine the $\kappa^{th}$ row of the matrix $\mathbf{A}^{*}$ where $\kappa=3,4,\ldots,\eta$ and positive integers $A_{k,\kappa}$ for $k=2,3,\ldots,\kappa-1$. By the formula $3A_{k,\kappa}-2A_{k-1,\kappa}=0$ for $\kappa=3,4,\ldots,\eta_{1}$ there are
$A_{k,\kappa}=2A_{k-1,\kappa}/3=2^{\eta-\kappa+k+\sigma(s)}3^{\kappa-k-1}\texttt{ for }k=2,3,\ldots,\kappa-1.$
So we obtain
(3.14) $A_{k,\kappa}=2^{\eta-\kappa+k+\sigma(s)}3^{\kappa-k-1}\texttt{ for }k=1,2,\ldots,\kappa-1.$
For $\xi=1,2,\ldots,s-1$ and $\kappa=\eta_{\xi}+1,\eta_{\xi}+2,\ldots,\eta_{\xi+1}$, there are
$A_{k,\kappa}=2A_{k-1,\kappa}/3=2^{\eta-\kappa+k+\sigma(\xi+1,s)}3^{\kappa-k-1}\texttt{ for }k=2,3,\ldots,\eta_{1}$
and for $\zeta=1,2,\ldots,\xi$ and $k=\eta_{\zeta}+1,\eta_{\zeta}+2,\ldots,\min(\eta_{\zeta+1},\kappa-1)$, if $k=\eta_{\zeta}+1$ then by the formula $3A_{k,\kappa}-2^{1+r_{\zeta}}A_{k-1,\kappa}=0$ there are
$A_{k,\kappa}=2^{1+r_{\zeta}}A_{k-1,\kappa}/3=2^{\eta-\kappa+k+\sigma(\xi+1,s)+\sigma(\zeta)}3^{\kappa-k-1}$
else by the formula $3A_{k,\kappa}-2A_{k-1,\kappa}=0$ there are
$A_{k,\kappa}=2A_{k-1,\kappa}/3=2^{\eta-\kappa+k+\sigma(\xi+1,s)+\sigma(\zeta)}3^{\kappa-k-1}$
for $k=\eta_{\zeta}+2,\eta_{\zeta}+3,\ldots,\min(\eta_{\zeta+1},\kappa-1)$. So for $\xi=1,2,\ldots,s-1$ and $\kappa=\eta_{\xi}+1,\eta_{\xi}+2,\ldots,\eta_{\xi+1}$ we obtain
(3.15) $A_{k,\kappa}=2^{\eta-\kappa+k+\sigma(\xi+1,s)+\sigma(\zeta)}3^{\kappa-k-1}$
for $\zeta=0,1,2,\ldots,\xi$ and $k=\eta_{\zeta}+1,\eta_{\zeta}+2,\ldots,\min(\eta_{\zeta+1},\kappa-1)$.
Thus by formulas 3.14-3.15 integers $A_{k,\kappa}$ are determined for $\kappa=2,3,\ldots,\eta$ and $k=1,2,\ldots,\kappa-1$.
Hence the matrix $\mathbf{A}^{*}$ is determined.
Let $A_{\kappa}=\sum_{j=1}^{\eta}A_{j,\kappa}$ for $\kappa=1,2,\ldots,\eta$. Then by formulas 3.12-3.15 there are
$A_{\kappa}=\sum_{k=1}^{\kappa-1}2^{\eta-\kappa+k+\sigma(s)}3^{\kappa-k-1}$
$\sum_{k=\kappa}^{\eta_{1}}2^{k-\kappa}3^{\eta-k+\kappa-1}+\sum_{\zeta=1}^{s-1}\sum_{k=\eta_{\zeta}+1}^{\eta_{\zeta+1}}2^{k-\kappa+\sigma(\zeta)}3^{\eta-k+\kappa-1}$
$2^{\eta-\kappa+1+\sigma(s)}3^{\kappa-1}[1-(2/3)^{\kappa-1}]+3^{\eta}[1-(2/3)^{\eta_{1}-\kappa+1}]$
$\sum_{\zeta=1}^{s-1}2^{\eta_{\zeta}+1-\kappa+\sigma(\zeta)}3^{\eta-\eta_{\zeta}+\kappa-1}[1-(2/3)^{\eta_{\zeta+1}-\eta_{\zeta}}]\texttt{ for }\kappa=1,2,\ldots,\eta_{1}$
and for $\xi=1,2,\ldots,s-1$ and $\kappa=\eta_{\xi}+1,\eta_{\xi}+2,\ldots,\min(\eta_{\xi+1},\eta)$
$A_{\kappa}=\sum_{\zeta=0}^{\xi}\sum_{k=\eta_{\zeta}+1}^{\min(\eta_{\zeta+1},\kappa-1)}2^{\eta-\kappa+k+\sigma(\xi+1,s)+\sigma(\zeta)}3^{\kappa-k-1}$
$\sum_{k=\kappa}^{\eta_{\xi+1}}2^{k-\kappa}3^{\eta-k+\kappa-1}+\sum_{\zeta=\xi+1}^{s-1}\sum_{k=\eta_{\zeta}+1}^{\eta_{\zeta+1}}2^{k-\kappa+\sigma(\xi+1,\zeta)}3^{\eta-k+\kappa-1}$
$\sum_{\zeta=0}^{\xi}2^{\eta-\kappa+\eta_{\zeta}+1+\sigma(\xi+1,s)+\sigma(\zeta)}3^{\kappa-\eta_{\zeta}-1}[1-(2/3)^{\min(\eta_{\zeta+1},\kappa-1)-\eta_{\zeta}}]$
$3^{\eta}[1-(2/3)^{\eta_{\xi+1}-\kappa+1}]+\sum_{\zeta=\xi+1}^{s-1}2^{\eta_{\zeta}+1-\kappa+\sigma(\xi+1,\zeta)}3^{\eta-\eta_{\zeta}+\kappa-1}[1-(2/3)^{\eta_{\zeta+1}-\eta_{\zeta}}].$
Particularly we have
$A_{1}=\sum_{\zeta=0}^{s-1}2^{\eta_{\zeta}+\sigma(\zeta)}3^{\eta-\eta_{\zeta}}[1-(2/3)^{\eta_{\zeta+1}-\eta_{\zeta}}].$
It can be rewritten as
$A_{1}=\sum_{\zeta=1}^{s}2^{\eta_{\zeta-1}+\sigma(\zeta-1)}3^{\eta-\eta_{\zeta}}(3^{\eta_{\zeta}-\eta_{\zeta-1}}-2^{\eta_{\zeta}-\eta_{\zeta-1}})$
or
$A_{1}=\sum_{\zeta=1}^{s-2}2^{\eta_{\zeta}+\sigma(\zeta)}3^{\eta-\eta_{\zeta}}-\sum_{\zeta=1}^{s-2}2^{\eta_{\zeta+1}+\sigma(\zeta)}3^{\eta-\eta_{\zeta+1}}$
$3^{\eta}-2^{\eta_{1}}3^{\eta-\eta_{1}}+2^{\eta_{s-1}+\sigma(s-1)}3^{\eta_{s}-\eta_{s-1}}-2^{\eta+\sigma(s-1)}$
$\sum_{\zeta=1}^{s}2^{\eta_{\zeta}+\sigma(\zeta-1)}3^{\eta-\eta_{\zeta}}(2^{\sigma(\zeta,\zeta)}-1)+Det\mathbf{A}.$